OpenStudy (anonymous):

Given F(x) = log8 x, find F(64).

OpenStudy (turingtest):

is this\[f(x)=\log_8x\]??

OpenStudy (anonymous):

yes just need to find f(64)

OpenStudy (turingtest):

no idea how to find\[F(64)=\log_8(64)\]??? Remember that 'log base b of x' is asking\[b^{(?)}=x\]so here we want to know\[8^{(?)}=64\]i.e. 8 to what power equals 64?

OpenStudy (anonymous):

im not sure i have a whole test like this

OpenStudy (turingtest):

Well you need to understand logarithms or else you are going to have trouble. what is \[8^2=?\]

OpenStudy (anonymous):

this is the problem. Given F(x) = log8 x, find F(64). 1/2 2 8

OpenStudy (turingtest):

can you please answer my question?\[8^2=?\]

OpenStudy (anonymous):

64

OpenStudy (turingtest):

right, and\[\log_8(64)\]is asking " 8 to what power equals 64?" and what power did we raise 8 to to get 64 ?

OpenStudy (anonymous):

2

OpenStudy (turingtest):

so there's you answer :-)

OpenStudy (turingtest):

\[\log_8(64)=2\]means\[8^2=64\]that is the meaning of a logarithm. In general\[\log_ba=x\]means\[b^x=a\]